OpenStudy (anonymous):

HELP PLZ! 3 capacitors of 20,10,15 uf respectively are connected on series across the 100v supply. find the voltage a cross the 10uf capacitors.

OpenStudy (lastdaywork):

Do you know the formula of net capacitance in series ??

OpenStudy (anonymous):

1/ct = 1/c1+1/c2+1/c3 or by reciprocal ct= c1xc2/c1+c2 ?

OpenStudy (lastdaywork):

Great! Then you can find the net charge by Q = CV As the charge on each capacitor will be same (why? ) you can find the voltage across any one of the capacitors using the same formula (Q = CV)

OpenStudy (anonymous):

i dont get it ... will u show me how?

OpenStudy (lastdaywork):

To find net charge (Q) substitute - C = net capacitance of series combination V = Voltage in the formula Q = CV Then, to find the the voltage across any capacitor (V') substitute - Q = net charge C' = capacitance of the concerned capacitor in the formula Q = C' * V'

OpenStudy (anonymous):

so 1st we need to get the value of C=1/20uf +1/10uf + 1/15uf ?

OpenStudy (lastdaywork):

Yea.. LHS would be 1/C

OpenStudy (anonymous):

so uf does have value?

OpenStudy (lastdaywork):

uf is a unit. A personal advice - "Before solving a question, convert all units to SI units and then forget them. Finally place the corresponding SI unit in the answer."